DP Gumby produces sleeveless sweaters. He can separate the demand function into
a domestic component:
QD = 32 – 0.4PD
and a foreign component
QF = 18 – 0.1PF
His total cost function is: C = 50 + 40Q
(a) On three separate diagrams sketch the demand curve in the domestic
market, the demand curve in the foreign market and the total demand curve.
1
Expert's answer
2019-04-02T10:18:56-0400

QD = 32 – 0.4PD or PD = 80 - 2.5QD,

QF = 18 – 0.1PF or PF = 180 - 10QF,

C = 50 + 40Q

(a) The demand curve in the domestic market is a graphical representation of the domestic demand function PD = 80 - 2.5QD, the demand curve in the foreign market is a graphical representation of the foreign demand function PF = 180 - 10QF and the total demand curve is a sum of domestic and foreign demand or P = 260 - 12.5Q..